How many milliliters of concentrated H2SO4 (18.0 M) are required to

prepare 750 ml of 3.00 M H2SO4 solution?

Solution:



1
Expert's answer
2021-02-01T03:56:21-0500

С 1 V 1 = С 2 V 2"\implies"2 = C 1 V 1 / V 2 = 750*3.00/18.0= 125 ml